OpenStudy (anonymous):

Why is the answer to this equation ln(4)?

OpenStudy (anonymous):

calculator?

OpenStudy (anonymous):

the problem is \[\int\limits_{e}^{4e} 1/x\]

OpenStudy (anonymous):

i know the antiderivative of 1/x is ln(x), correct? but idk where to go from there

OpenStudy (turingtest):

evaluate it from those bounds is all ln(4e)-ln(e)=ln(4)

OpenStudy (anonymous):

that's what I thought. I can see with a calculator that ln(4e)-ln(e)=ln(4), but how to I figure that out without a calculator?

OpenStudy (turingtest):

\[\int_{e}^{4e}\frac{1}{x}dx=\ln x|_{e}^{4e}=\ln(4e)-\ln e=\ln4+\ln e-\ln e=\ln4\]remember your log rules!\[\log(ab)=\log a+\log b\]

OpenStudy (anonymous):

got it! thank you!

OpenStudy (turingtest):

welcome :)